If 3 years ago sara was 2 times john's age and 9 years from now she is 1.5 times john's age, how old are they now.
Answer by WannabeCAgirl83(35)

Sara = s
John = j

1) If 3 years ago sara was 2 times john's age...
1) s - 3 = 2j

2) and 9 years from now she is 1.5 times john's age...
2) s - 9 = 1.5j

...how old are they now?

Step 1
Isolating s in equation 1):
s - 3 = 2j | +3
s = 2j + 3

Step 2
Substituting s with 2j+3 in equation 2):
s - 9 = 1.5j
2j + 3 - 9 = 1.5j
2j - 6 = 1.5j | -2j
-6 = -0.5j | : -0.5
12 = j → John is 12 years old.

Step 3
Substituting j with 12 in equation 1):
s - 3 = 2j
s - 3 = 2 ∙ 12
s - 3 = 24 | +3
s = 27 → Sara is 27 years old.

Step 4
Validating solutions by substituting s with 27 and j with 12 to check if both sides of the equations are really equal:
1) s - 3 = 2j
1) 27 - 3 = 2 ∙ 12
1) 24 = 24

2) s - 9 = 1.5j
2) 27 - 9 = 1.5 ∙ 12
2) 18 = 18

As both sides of each equation are equal s = 27 and j = 12 are valid solutions.